# A Two-Stage GPU Kernel Tuner Combining Semantic Refactoring and Search-Based Optimization

### v2

@@
Abstract:  To address these challenges, we propose a two-level GPU code tuner that combines semantic refactoring with search-based optimization. Building on an agentic iterative loop, our method introduces template-based rewriting that explicitly refactors candidate kernels into parameterizable templates, and then performs search-based auto-tuning over the template parameters to obtain more stable and higher-quality performance improvements. We extract real-world kernels from SGLang as evaluation targets. The proposed agentic tuner completes templating, testing, analysis, and planning within the iterative loop, and conducts parameter search under hardware constraints guided by profiling signals. Experiments show that, compared with SGLang’s native kernel implementations, our method achieves over 3× speedup on a set of real kernels. Compared with agent-only direct rewriting, templating plus search significantly reduces iteration randomness, makes the optimization process more interpretable, and approaches high-performance configurations in a more systematic manner. The approach can be further extended to backends such as OpenCL and HIP, providing automated performance optimization for real serving workloads.
Keywords: search-based optimization, CUDA, GPU, auto-tuning, agents

### v3

@@
Abstract:  GPU code optimization is a key bottleneck for high-performance computing (HPC) and for large-model training and inference. Expert-written kernels require semantics-level refactoring and fine-grained parameter tuning, which is costly, slow, and hard to scale. Although LLM-based agents have reduced the barrier to kernel generation and optimization, most existing approaches directly rewrite code, leaving critical performance choices implicit and lacking controllable parameterized interfaces and systematic search; consequently, speedups can be unstable and the optimization process is difficult to interpret and reproduce. We propose a two-level GPU code tuner that couples semantic refactoring with search-based optimization: within an agentic iterative feedback loop, candidate kernels are explicitly rewritten into parameterizable templates, and template parameters are tuned via guided search under hardware constraints using profiling signals. We evaluate on real-world kernels extracted from SGLang, where our method achieves over 3× speedup versus native SGLang implementations on a set of kernels. Compared with agent-only direct rewriting, templating plus search reduces iteration randomness, improves interpretability and reproducibility, and more systematically approaches high-performance configurations. The framework can be extended to other backends such as OpenCL and HIP to provide automated optimization for practical service workloads.